Update #3: The Band 12 detector was switched on Apr 08, 2021 at 1800Z.  Following a performance evaluation engineers determined that this action resolved the horizontal striping anomaly on Band 12.

Update #2: GOES-17 engineers switched detectors to correct the horizontal striping problem on Band 12.  Band 12 performance is being monitored.

Update #1: This anomaly is due to horizontal striping on Band 12.
Topic: GOES-17 ABI L1b Band 12Cloud and Moisture Band 12Sectorized Cloud and Moisture Band 12 and derived products delivered to AWIPSGRBPDA
Date/Time Issued: Apr 08, 2021 1935Z
Product or Data Impacted: GOES-17 ABI L1b Band 12Cloud and Moisture Band 12Sectorized Cloud and Moisture Band 12 and derived products All Scenes delivered to AWIPSGRBPDA.
Date/Time of Initial Impact: Apr 02, 2021 1613Z
Date/Time of Expected End: TBD
Length of the Outage/Event: 
Details/Specifics of Change: GOES-17 ABI L1b Band 12Cloud and Moisture Band 12Sectorized Cloud and Moisture Band 12 and derived products All Scenes delivered to AWIPSGRBPDA. Due to an anomaly. 
contact information for further information: ESPC Operations at ESPCOperations@noaa.gov or 301-817-3880 24x7 or ross.cox@noaa.gov during normal business hours.

Web Site(s) for applicable information: N/A